Suppose that a consumer consumes two goods X and Y and derives utility according the following utility function where U = 25X2/5Y 3/5 where α = 2/5 and β = 3/5 a. If Px is the price of good X and Py is the price of good Y and the consumer’s income is M. Derive the demand functions for the two goods X and Y b. If Px is shs 15 and Py is shs 10 and the consumer has shs.800 to spend on the two goods what are the optimal quantities of X and Y that maximize the consumer’s utility? c. Using the information in b above show that the values of α and β represent the proportion of the consumer’s income spent on good X and good Y respectively


1
Expert's answer
2022-02-01T06:55:35-0500

"U= 25X^\frac{2}{5}Y^\frac{3}{5}"

"\alpha= \frac{2}{5}"

"\beta= \frac{3}{5}"

a) Demand functions

Budget = "P_xX+ P_yY= m"

"\frac{Mu_x}{Mu_y}= \frac{p_x}{P_y}"

"Mu_x= 10X^\frac{-3}{5}Y\frac{3}{5}"


"Mu_y= 15X^\frac{2}{5}Y\frac{-2}{5}"

"\frac{10X^\frac{-3}{5}Y\frac{3}{5}}{15X^\frac{2}{5}Y\frac{-2}{5}}= \frac{P_x}{P_y}"


"\frac{10Y}{15X}= \frac{P_x}{P_y}"


"15P_xX=10P_yY"


"X= \frac{0.67P_yY}{P_x}"

"Y= \frac{1.5P_xX}{P_y}"


Plug into the budget Equation

"P_xX+P_yY=m"


"P_x(\frac{0.67P_yY}{P_x})+P_yY=m"


"0.67P_yY+P_yY=m"


"\frac{2}{3}P_y+P_yY=m"

"5P_y= 3m"


"Y^*= \frac{3m}{5P_y}"

"P_xX+P_y(\frac{1.5P_xX}{P_y})=m"

"2.5P_xX=m"


"X^*= \frac{m}{2.5P_x}"


"= \frac{m} {\frac{5}{2}P_x}"


"X^*= \frac{2m}{5P_x}"

b) Optimal Quantities


"Y^*= \frac{3m}{5P_y}= \frac{3\times800}{5\times10}"


"=\frac{2400}{50}= 48"


"X^*= \frac{m}{2.5P_x}= \frac{800}{2.5\times 5}= 64"


Amount spend on good X="64\times5=320"

Amount spend on good Y="48\times 10= 480"

Proportion on X= "\frac{320}{800}=0.4=\alpha"


Proportion on Y= "\frac{480}{800}=0.6=\beta"